A fortress has been built on top of a cliff. You measure the angle of elevation of the top of the cliff from a point 200 ft horizontally away from the base of the cliff to be 30' and the angle of elevation of the top of the fortress wall from the same point to be 45'. What is the height of the fortress wall?
Answer by mananth(16946)

Angle of elevation of cliff base = 30/60 deg ( 60 ' = 1 deg)
distance on ground = 200 feet
so height of cliff = 200*Tan(1/2)
..
Angle of elevation of wall top = 45/60 deg
distance on ground = 200 feet
so height of cliff+ wall = 200*Tan(3/4)

200*tan(3/4)-200*tan(1/2) = height of wall
= 77.05 feet